What is the result of the following expression

30 + 10.0 * 5 - 50 % 5

in C++ please
Answer #1

In C++, for arithmetic operator *, /, % have higher precedence than +, - And they are evaluated from left to right.

Here, 30 + 10.0 * 5 - 50 % 5

So first 10.0*5 will be evaluated. 10.0*5 = 50.0

Now, the expression has become: 30 + 50.0 - 50 % 5

Then, 50%5 will be evaluated. 50%5 = 0

Now, the expression has become: 30 + 50.0 - 0

Then, 30 + 50.0 will be evaluated. 30 + 50.0 = 80

Now, the expression has become: 80 - 0

So, finally, 80 - 0 will be evaluated. 80 - 0 = 80

Therefore, reuslt = 80